AS1805(1L) AS1805(1L) NAME as1805,as1802−crossassemblersformicrocomputers SYNOPSIS as1805[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input as1802[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas1805,as1802command(s)assemblestheinputfileintoatextoutputfilerepresentingthe programmemoryforamicrocomputer. Options -pcpuOverridetheinstructionsetselection. ValidValuesforCpu 1802,1804,1805,1806 -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas1805Frankenstein Assembler(fileas1805.doc) NOTES Thereisonlyoneinputfile. Theprogramcanselectwhichsubsetofinstructionsisaccepted.Theprogramfirstusesthename withwhichitisinvoked.Thisisscannedforasubstringwhichindicateswhichsettouse.The-p optionsoverridesthisselctionbyperformingthesamesubstringsearch. Thereshouldonlybeoneexecutablefilewithallofitsnameslinkedtoit. 1